# Shanghai 200433, China

### v1

@@ Appendix B Asymptotic Analysis of Fourier Integral
$I(V)\sim\int_{c-\epsilon}^{c+\epsilon}\left[h(c)+h^{\prime}(c)(\mu-c)+\frac{1}{2}h^{\prime\prime}(c)(\mu-c)^{2}\right]\\ \times\exp\left\{V\left[g(c)+\frac{1}{2}(\mu-c)^{2}g^{\prime\prime}(c)+\frac{1}{6}(\mu-c)^{3}g^{\prime\prime\prime}(c)\right.\right.+\left.\left.\frac{1}{24}(\mu-c)^{4}g^{(4)}(c)\right]\right\},$ (105)
@@
$I(V)\sim\int_{-\infty}^{+\infty}\mathrm{e}^{Vg(c)+V(\mu-c)^{2}g^{\prime\prime}(c)/2}\\ \times\left\{h(c)+\frac{(\mu-c)^{2}}{2}h^{\prime\prime}(c)+(\mu-c)^{4}\left[\frac{1}{24}Vh(c)g^{(4)}(c)+\frac{1}{6}Vh^{\prime}(c)g^{\prime\prime\prime}(c)\right]\right.\\ \left.+\frac{1}{72}(\mu-c)^{6}V^{2}h(c)[g^{\prime\prime\prime}(c)]^{2}\right\},\quad V\rightarrow+\infty.$ (107)
@@
$I(V)=\frac{\mathrm{e}^{Vg}}{2\pi}\frac{1}{\sqrt{V(-g^{\prime\prime})}}\left(h+\frac{-\frac{h^{\prime\prime}}{2g^{\prime\prime}}+\frac{g^{(3)}h^{\prime}}{2(g^{\prime\prime})^{2}}-\frac{5h\left(g^{(3)}\right)^{2}}{24(g^{\prime\prime})^{3}}+\frac{hg^{(4)}}{8(g^{\prime\prime})^{2}}}{V}\right).$ (108)
Here, $g^{(k)}$ denotes the $k$-th derivative with respect to $\mu$, and both the function $h$ and the function $g$, along with their derivatives, are evaluated at $\mu=c$. The appearance of $-g^{\prime\prime}(c)$ in the denominator under the square root arises because $c$ is the point where $g$ attains its maximum, meaning its second derivative is negative, $g^{\prime\prime}(c)<0$.

### v2

@@ Appendix C Asymptotic Analysis of Fourier Integral
$I(V)\sim\int_{c-\epsilon}^{c+\epsilon}\left[h(c)+h^{\prime}(c)(\mu-c)+\frac{1}{2}h^{\prime\prime}(c)(\mu-c)^{2}\right]\\ \times\exp\left\{V\left[g(c)+\frac{1}{2}(\mu-c)^{2}g^{\prime\prime}(c)+\frac{1}{6}(\mu-c)^{3}g^{\prime\prime\prime}(c)\right.\right.+\left.\left.\frac{1}{24}(\mu-c)^{4}g^{(4)}(c)\right]\right\}\mathrm{d}\mu,$ (119)
@@
$I(V)\sim\int_{-\infty}^{+\infty}\mathrm{e}^{Vg(c)+V(\mu-c)^{2}g^{\prime\prime}(c)/2}\\ \times\left\{h(c)+\frac{(\mu-c)^{2}}{2}h^{\prime\prime}(c)+(\mu-c)^{4}\left[\frac{1}{24}Vh(c)g^{(4)}(c)+\frac{1}{6}Vh^{\prime}(c)g^{\prime\prime\prime}(c)\right]\right.\\ \left.+\frac{1}{72}(\mu-c)^{6}V^{2}h(c)[g^{\prime\prime\prime}(c)]^{2}\right\}\mathrm{d}\mu,\quad V\rightarrow+\infty.$ (121)
@@
$I(V)=\mathrm{e}^{Vg}\frac{\sqrt{2\pi}}{\sqrt{V(-g^{\prime\prime})}}\left(h+\frac{-\frac{h^{\prime\prime}}{2g^{\prime\prime}}+\frac{g^{(3)}h^{\prime}}{2(g^{\prime\prime})^{2}}-\frac{5h\left(g^{(3)}\right)^{2}}{24(g^{\prime\prime})^{3}}+\frac{hg^{(4)}}{8(g^{\prime\prime})^{2}}}{V}\right).$ (122)
Here, $g^{(k)}$ denotes the $k$-th derivative with respect to $\mu$, and both the function $h$ and the function $g$, along with their derivatives, are evaluated at $\mu=c$. The appearance of $g^{\prime\prime}(c)$ in the denominator under the square root arises because $c$ is the point where $g$ attains its maximum, meaning its second derivative is negative, $g^{\prime\prime}(c)<0$.